# A Flexible, Equivariant Framework for Subgraph GNNs via Graph Products and Graph Coarsening

### v1

@@ Abstract
Subgraph Graph Neural Networks (Subgraph GNNs) enhance the expressivity of message-passing GNNs by representing graphs as sets of subgraphs. They have shown impressive performance on several tasks, but their complexity limits applications to larger graphs. Previous approaches suggested processing only subsets of subgraphs, selected either randomly or via learnable sampling. However, they make suboptimal subgraph selections or can only cope with very small subset sizes, inevitably incurring performance degradation. This paper introduces a new Subgraph GNNs framework to address these issues. We employ a graph coarsening function to cluster nodes into super-nodes with induced connectivity. The product between the coarsened and the original graph reveals an implicit structure whereby subgraphs are associated with specific sets of nodes. By running generalized message-passing on such graph product, our method effectively implements an efficient, yet powerful Subgraph GNN. Controlling the coarsening function enables meaningful selection of any number of subgraphs while, contrary to previous methods, being fully compatible with standard training techniques. Notably, we discover that the resulting node feature tensor exhibits new, unexplored permutation symmetries. We leverage this structure, characterize the associated linear equivariant layers and incorporate them into the layers of our Subgraph GNN architecture. Extensive experiments on multiple graph learning benchmarks demonstrate that our method is significantly more flexible than previous approaches, as it can seamlessly handle any number of subgraphs, while consistently outperforming baseline approaches.
@@ 1 Introduction
Subgraph GNNs [4, 12, 40, 8, 28, 30, 39, 3] have recently emerged as a promising direction in graph neural network research, addressing the expressiveness limitations of Message Passing Neural Networks (MPNNs) [24, 36, 25]. In essence, a Subgraph GNN operates on a graph by transforming it into a collection of subgraphs, generated based on a specific selection policy. For example, this could involve removing a single node from the original graph or simply marking a node without changing the graph’s original connectivity [27]. The model then processes these subgraphs using an equivariant architecture, aggregates the derived representations, and makes graph- or node-level predictions. The growing popularity of Subgraph GNNs stems not only from their enhanced expressive capabilities over MPNNs but also from their impressive empirical results, as notably demonstrated on well-known molecular benchmarks [39, 12, 3].
Unfortunately, Subgraph GNNs are hindered by substantial computational costs as they necessitate message-passing operations across all subgraphs within the bag. Typically, the number of subgraphs is the number of nodes in the graph, $n$, resulting in a time complexity scaling quadratically ($\mathcal{O}(n^{2})$) for bounded degree graphs, in contrast to a standard MPNN, whose time complexity is linear in the number of nodes for such graphs. This significant computational burden becomes impractical in large graphs, limiting the applicability of Subgraph GNNs to important applications and widely used datasets. To overcome this challenge, various studies have explored methodologies that process only a subset of subgraphs from the bag. These methods range from simple random sampling techniques [8, 4, 41, 3] to more advanced strategies that learn to select the most relevant subset of the bag to process [5, 20, 30]. However, while random sampling of subgraphs yields subpar performance, more sophisticated learnable selection strategies also have significant limitations. Primarily, they rely on discrete sampling during training, which can complicate the training process, as evidenced by the high number of epochs required to train them  [20, 5, 30]. As a result, these methods often allow only a very small bag size, which only yields modest performance improvements compared to random sampling and standard MPNNs.
@@
The goal of this paper is to devise a Subgraph GNN architecture that can flexibly generate and process variable-sized subgraph policies, and deliver strong experimental results while sidestepping intricate and lengthy training protocols. Specifically, our approach aims to overcome the common limitation of restricting usage to a very small set of subgraphs.
@@
Our proposed method builds upon and extends an observation made in [3], which draws an analogy between using Subgraph GNNs and performing message-passing operations over a larger “product graph”. Specifically, it was shown that when considering the maximally expressive Subgraph GNN suggested by [39], the bag of subgraphs and its update rules can be obtained by transforming a graph through the graph cartesian product of the original graph with itself, i.e. $G\Box G$, and then processing this resulting graph using a standard MPNN. In our approach, we propose to modify the first term of the product and replace it with a coarsened version of the original graph, denoted $\mathcal{T}(G)$, obtained by mapping nodes to super-nodes (e.g., by applying graph clustering, see Figure 1(left)), making the resulting product graph $\mathcal{T}(G)\Box G$ significantly smaller. This construction is illustrated in Figure 1(right). This process effectively associates each subgraph with a set of nodes, so that, by controlling the coarsening function, it allows for both flexible bag sizes and a simple, meaningful selection of the subgraphs.
@@
While performing message passing on $\mathcal{T}(G)\Box G$ serves as the core update rule in our architecture, we augment our message passing operations with another set of operations derived from the symmetry structure of the resulting node feature tensor, which we call symmetry-based updates. Specifically, the node feature tensor in our case is indexed by pairs $(S,i)$ where $S$ is a super-node and $i$ is an original node. Accordingly, $\mathcal{X}$ is a ${2^{n}\times n\times d}$ (very) sparse tensor, where $2^{n}$ is the number of all subsets of the vertex set, and $d$ is the feature dimension. Interestingly, we find that this object adheres to a specific set of symmetries, which, to the best of our knowledge, is yet unstudied in the context of machine learning: applying permutation $\sigma\in S_{n}$ to the nodes in $S$ and to $i$ results in an equivalent representation of our node feature tensor. We formally define the symmetries of this object and characterize all the affine equivariant operations in this space. We incorporate these operations into our message-passing by encoding the parameter-sharing schemes [31] as additional edge features, and by deriving novel provably powerful node marking techniques for our setup. These additional update rules significantly improve experimental results. We note, that our symmetry analysis may be useful for processing bags derived from other high-order generation policies [30, 20] by treating tuples of nodes as sets.
The flexibility and effectiveness of our entire framework are illustrated in Figure 2, in which we performed a thorough experimental study on the popular Zinc-12k dataset [32]. Our method demonstrates a significant performance boost over baseline models in the small bag setting (the one they are designed for), while achieving results that compare favourably to state-of-the-art Subgraph GNNs in the full bag setting. Additionally, we can obtain results in-between these two regimes.
@@
The main contributions of this paper are: (1) the development of a novel, flexible Subgraph GNN framework that enables meaningful construction and processing of bags of subgraphs of any size; (2) a characterization of all affine invariant/equivariant layers defined on our node feature tensors; (3) a theoretical analysis of our method, demonstrating the expressivity benefits of our approach; (4) a comprehensive experimental evaluation demonstrating the advantages of the new approach across both small and large bag sizes, achieving state-of-the-art results, often by a significant margin.
@@ 2 Related work
Subgraph GNNs. Subgraph GNNs [40, 8, 28, 4, 41, 27, 12, 30, 17, 39, 3] represent a graph as a collection of subgraphs, obtained by a predefined generation policy. For example, each subgraph can be generated by marking exactly one node in the original graph (see inset  [^1] The Figure was taken with permission from [3]) – an approach commonly referred to as node marking [27]; this marked node is considered the root node in his subgraph. Several recent papers focused on scaling these methods to larger graphs, starting with basic random selection of subgraphs from the bag, and extending beyond with more sophisticated techniques that aim to learn how to select subgraphs. To elaborate, [5] introduced Policy-Learn (PL), an approach based on two models, where the first model predicts a distribution over the nodes of the original graph, and the second model processes bags of subgraphs sampled from this distribution. MAG-GNN [20] employs a similar approach utilizing reinforcement learning. Similarly to our approach, this method permits high-order policies by associating subgraphs with tuples rather than individual nodes, allowing for the marking of several nodes within a subgraph. However, as mentioned before, these approaches involve discrete sampling while training, making them very hard to train (1000-4000 epochs vs. $\sim$400 epochs of state-of-the-art methods [3, 39] on the Zinc-12k dataset), and limiting their usage to very small bags. Finally, we mention another high-order method, OSAN, introduced by [30], which learns a distribution over tuples that represent subgraphs with multiple node markings. In the context of these works, our subgraph selection procedure can be seen as a variant of high-order policies, where instead of using ordered tuples, our subgraphs are indexed by subsets of indices. In contrast to these previous approaches, we suggest a simple and effective way to select subgraphs and also show how to leverage the resulting symmetry structure to augment our message-passing operations.
Symmetries in graph learning. Many previous works have analyzed and utilized the symmetry structure that arises from graph learning setups [22, 23, 18, 2]. Specifically relevant to our paper is the work of [22] that characterized basic equivariant linear layers for graphs, the work of [1] that characterizes equivariant maps for many other types of incidence tensors that arise in graph learning, and the recent works [4, 12] that leveraged group symmetries for designing Subgraph GNNs in a principled way.
@@ 3 Preliminaries
Notation. Let $\mathcal{G}$ be a family of undirected graphs, and consider a graph $G=(V,E)$ within this family. The adjacency matrix $A\in\mathbb{R}^{n\times n}$ defines the connectivity of the graph [^2] Edge features are omitted for simplicity., while the feature matrix $X\in\mathbb{R}^{n\times d}$ represents the node features. Here, $V$ and $E$ represent the sets of nodes and edges, respectively, with $|V|=n$ indicating the number of nodes. We use the notation $v_{1}\sim_{A}v_{2}$ to denote that $v_{1}$ and $v_{2}$ are neighboring nodes according to the adjacency $A$. Additionally, we define $[n]\triangleq\{1,2,\ldots n\}$, and $\mathcal{P}([n])$ as the power set of $[n]$.
Subgraph GNNs as graph products. In a recent work, [3] demonstrated that various types of update rules used by current Subgraph GNNs can be simulated by employing a Cartesian graph product between the original graph and another graph, and running standard message passing over that newly constructed product graph. Formally, the cartesian product of two graphs $G_{1}$ ($n_{1}$ nodes) and $G_{2}$ ($n_{2}$ nodes), denoted by $G_{1}\square G_{2}$, forms a graph with vertex set $V(G_{1})\times V(G_{2})$. Two vertices $(u_{1},u_{2})$ and $(v_{1},v_{2})$ are adjacent if either $u_{1}=v_{1}$ and $u_{2}$ is adjacent to $v_{2}$ in $G_{2}$, or $u_{2}=v_{2}$ and $u_{1}$ is adjacent to $v_{1}$ in $G_{1}$. We denote by $\mathcal{A}\in\mathbb{R}^{n_{1}\cdot n_{2}\times n_{1}\cdot n_{2}}$ and $\mathcal{X}\in\mathbb{R}^{n_{1}\cdot n_{2}\times d}$ the adjacency and node feature matrices of the product graph; in general, we use calligraphic letters to denote the adjacency and feature matrices of product graphs, while capital English letters are used for those of the original graphs. In particular, for the graph cartesian product, $G_{1}\square G_{2}$, the following holds:
@@
For a detailed definition of the cartesian product of graphs, please refer to Definition A.1. As a concrete example for the analogy between Subgraph GNNs and the Cartesian product of graphs, we refer to a result by [3], which states that the maximally expressive Subgraph GNN architecture GNN-SSWL$+$ [39], can be simulated by an MPNN on the Cartesian product of the original graph with itself, denoted as $G\square G$. As we shall see, our framework utilizes a cartesian product of the original graph and a coarsened version of it, as illustrated in Figure 1 (right).
Equivariance. A function $L:V\rightarrow W$ is called equivariant if it commutes with the group action. More formally, given a group element, $g\in\mathbb{G}$, the function $L$ should satisfy $L(g\cdot v)=g\cdot L(v)$ for all $v\in V$ and $g\in\mathbb{G}$. Invariance, a related concept, occurs when $L$ satisfies $L(g\cdot v)=L(v)$.
@@ 4 Coarsening-based Subgraph GNN
Overview. This section introduces the Coarsening-based Subgraph GNN (CS-GNN) framework. The main idea is to select and process subgraphs in a principled and flexible manner through the following approach: (1) coarsen the original graph via a coarsening function, $\mathcal{T}$ – see Figure 1(left); (2) Obtain the product graph – Figure 1(right) defined by the combination of two adjacencies, $\mathcal{A}_{\mathcal{T}(G)}$ (red edges), $\mathcal{A}_{G}$ (grey edges), which arise from the graph Cartesian product operation (details follow); (3) leveraging the symmetry of this product graph to develop symmetry-based updates, described by $\mathcal{A}_{\text{Equiv}}$ (this part is not visualized in Figure 1). The general update of our suggested layer takes the following form [^3] Edge features are also allowed but are omitted here for simplicity.,
@@
where the superscript t indicates the layer index. In what follows, we further elaborate on these three steps (in Sections 4.1, 4.1 and 4.2). We note, that each connectivity in Equation 2 is processed using a different MPNN, for more specific implementation details, we refer to Appendix F.
4.1 Product graph definition
As mentioned before, a maximally expressive Subgraph GNN can be realized via the Cartesian product of the original graph with itself $G\square G$. In this work, we extend this concept by allowing the left operand in the product to be the coarsened version of $G$, denoted as $\mathcal{T}(G)$, as defined next. This idea is illustrated in Figure 1.

### v4

@@ Abstract
Subgraph GNNs enhance message-passing GNNs expressivity by representing graphs as sets of subgraphs, demonstrating impressive performance across various tasks. However, their scalability is hindered by the need to process large numbers of subgraphs. While previous approaches attempted to generate smaller subsets of subgraphs through random or learnable sampling, these methods often yielded suboptimal selections or were limited to small subset sizes, ultimately compromising their effectiveness. This paper introduces a new Subgraph GNN framework to address these issues. Our approach diverges from most previous methods by associating subgraphs with node clusters rather than with individual nodes. We show that the resulting collection of subgraphs can be viewed as the product of coarsened and original graphs, unveiling a new connectivity structure on which we perform generalized message passing.
Crucially, controlling the coarsening function enables meaningful selection of any number of subgraphs. In addition, we reveal novel permutation symmetries in the resulting node feature tensor, characterize associated linear equivariant layers, and integrate them into our Subgraph GNN. We also introduce novel node marking strategies and provide a theoretical analysis of their expressive power and other key aspects of our approach. Extensive experiments on multiple graph learning benchmarks demonstrate that our method is significantly more flexible than previous approaches, as it can seamlessly handle any number of subgraphs, while consistently outperforming baseline approaches. Our code is available at https://github.com/BarSGuy/Efficient-Subgraph-GNNs.
@@ 1 Introduction
Subgraph GNNs [4, 12, 40, 8, 28, 30, 39, 3] have recently emerged as a promising direction in graph neural network research, addressing the expressiveness limitations of Message Passing Neural Networks (MPNNs) [24, 36, 25]. In essence, a Subgraph GNN operates on a graph by transforming it into a collection of subgraphs, generated based on a specific selection policy. Examples of such policies include removing a single node from the original graph or simply marking a node without changing the graph’s original connectivity [27]. The model then processes these subgraphs using an equivariant architecture, aggregates the derived representations, and makes graph- or node-level predictions. The growing popularity of Subgraph GNNs stems not only from their enhanced expressive capabilities over MPNNs but also from their impressive empirical results, as notably demonstrated on well-known molecular benchmarks [39, 12, 3].
Unfortunately, Subgraph GNNs are hindered by substantial computational costs as they necessitate message-passing operations across all subgraphs within the bag. Typically, the number of subgraphs is the number of nodes in the graph, $n$— for bounded degree graphs, this results in a time complexity scaling quadratically ($\mathcal{O}(n^{2})$), in contrast to the linear complexity of a standard MPNN. This significant computational burden makes Subgraph GNNs impractical for large graphs, hindering their applicability to important tasks and widely used datasets. To overcome this challenge, various studies have explored methodologies that process only a subset of subgraphs from the bag. These methods range from simple random sampling techniques [8, 4, 41, 3] to more advanced strategies that learn to select the most relevant subset of the bag to process [5, 20, 30]. However, while random sampling of subgraphs yields subpar performance, more sophisticated learnable selection strategies also have significant limitations. Primarily, they rely on training-time discrete sampling which complicates the optimization process, as evidenced by the high number of epochs required to train them  [20, 5, 30]. As a result, these methods often allow only a very small bag size, yielding only modest performance improvements compared to random sampling and standard MPNNs.
@@
The goal of this paper is to devise a Subgraph GNN architecture that can flexibly generate and process variable-sized bags, and deliver strong experimental results while sidestepping intricate and lengthy training protocols. Specifically, our approach aims to overcome the common limitation of restricting usage to a very small set of subgraphs.
@@
Our proposed method builds upon and extends an observation made by Bar-Shalom et al. [3], who draw an analogy between using Subgraph GNNs and performing message-passing operations over a larger “product graph”. Specifically, it was shown that when considering the maximally expressive (node-based) Subgraph GNN suggested by [39] [^1] The architecture suggested in [39] was shown to be at least as expressive as all previously studied node-based Subgraph GNNs, the bag of subgraphs and its update rules can be obtained by transforming a graph through the graph cartesian product of the original graph with itself, i.e., $G\Box G$, and then processing the resulting graph using a standard MPNN. In our approach, we propose to modify the first term of the product and replace it with a coarsened version of the original graph, denoted $\mathcal{T}(G)$, obtained by mapping nodes to super-nodes (e.g., by applying graph clustering, see Figure 1(left)), making the resulting product graph $\mathcal{T}(G)\Box G$ significantly smaller. This construction is illustrated in Figure 1(right). This process effectively associates each subgraph – a row in Figure 1(right) – with a set of nodes produced by the coarsening function $\mathcal{T}$. Different choices of $\mathcal{T}$ allow for both flexible bag sizes and a simple, meaningful selection of the subgraphs.
While performing message passing on $\mathcal{T}(G)\Box G$ serves as the core update rule in our architecture, we augment our message passing operations with another set of operations derived from the symmetry structure of the resulting node feature tensor, which we call symmetry-based updates. Specifically, our node feature tensor is indexed by pairs $(S,v)$ where $S$ is a super-node and $v$ is an original node. Accordingly, $\mathcal{X}$ is a ${T\times n\times d}$ tensor, where $d$ is the feature dimension, and $T$ is the number of super-nodes (a constant hyper-parameter). As super-nodes are sets of nodes, $\mathcal{X}$ can also be viewed as a (very) sparse ${2^{n}\times n\times d}$ tensor where $2^{n}$ is the number of all subsets of the vertex set. Since the symmetric group $S_{n}$ acts naturally on this representation, we use it to develop symmetry based updates.
Interestingly, we find that this node feature tensor, $\mathcal{X}$, adheres to a specific set of symmetries, which, to the best of our knowledge, is yet unstudied in the context of machine learning: applying a permutation $\sigma\in S_{n}$ to the nodes in $S$ and to $v$ results in an equivalent representation of our node feature tensor. We formally define the symmetries of this object and characterize all the affine equivariant operations in this space. We incorporate these operations into our message-passing by encoding the parameter-sharing schemes [31] as additional edge features. These additional update rules significantly improve experimental results. We note that our symmetry analysis may be useful for processing bags derived from other high-order generation policies [30, 20] by treating tuples of nodes as sets.
Inspired by these symmetries and traditional binary-based [4] and shortest path-based [39] node-marking strategies, we propose four natural marking strategies for our framework. Interestingly, unlike the full-bag scenario, they vary in expressiveness, with the shortest path-based technique being the most expressive.
@@
The flexibility and effectiveness of our full framework are illustrated in Figure 2, depicting detailed experimental results on the popular Zinc-12k dataset [32]. Our method demonstrates a significant performance boost over baseline models in the small bag setting (for which they are designed), while achieving results that compare favourably to state-of-the-art Subgraph GNNs in the full bag setting. Additionally, we can obtain results in-between these two regimes.
@@
The main contributions of this paper are: (1) the development of a novel, flexible Subgraph GNN framework that enables meaningful construction and processing of bags of subgraphs of any size; (2) a characterization of all affine invariant/equivariant layers defined on our node feature tensors; (3) a theoretical analysis of our framework, including the expressivity benefits of our node-marking strategy; and (4) a comprehensive experimental evaluation demonstrating the advantages of the new approach across both small and large bag sizes, achieving state-of-the-art results, often by a significant margin.
@@ 2 Related work
Subgraph GNNs. Subgraph GNNs [40, 8, 28, 4, 41, 27, 12, 30, 17, 39, 3] represent a graph as a collection of subgraphs, obtained by a predefined generation policy. For example, each subgraph can be generated by marking exactly one node in the original graph (see inset  [^2] The Figure was taken with permission from [3]) – an approach commonly referred to as node marking [27]; this marked node is considered the root node in its subgraph. Several recent papers focused on scaling these methods to larger graphs, starting with basic random selection of subgraphs from the bag, and extending beyond with more sophisticated techniques that aim to learn how to select subgraphs. To elaborate, [5] introduced Policy-Learn (PL), an approach based on two models, where the first model predicts a distribution over the nodes of the original graph, and the second model processes bags of subgraphs sampled from this distribution. MAG-GNN [20] employs a similar approach utilizing Reinforcement Learning. Similarly to our approach, this method permits high-order policies by associating subgraphs with tuples rather than individual nodes, allowing for the marking of several nodes within a subgraph. However, as mentioned before, these approaches involve discrete sampling while training, making them very hard to train (1000-4000 epochs vs. $\sim$400 epochs of state-of-the-art methods [3, 39] on the Zinc-12k dataset), and limiting their usage to very small bags. Finally, we mention another high-order method, OSAN, introduced by [30], which learns a distribution over tuples that represent subgraphs with multiple node markings. In contrast to these previous approaches, we suggest a simpler and more effective way to select subgraphs and also show how to leverage the resulting symmetry structure to augment our message-passing operations.
Symmetries in graph learning. Many previous works have analyzed and utilized the symmetry structure that arises from graph learning setups [22, 23, 18, 2]. Specifically relevant to our paper is the work of [22] that characterized basic equivariant linear layers for graphs, the work of [1] that characterizes equivariant maps for many other types of incidence tensors that arise in graph learning, and the works [4, 12] that leveraged group symmetries for designing Subgraph GNNs in a principled way.
@@ 3 Preliminaries
Notation. Let $\mathcal{G}$ be a family of undirected graphs, and consider a graph $G=(V,E)$ within this family. The adjacency matrix $A\in\mathbb{R}^{n\times n}$ defines the connectivity of the graph [^3] Edge features are also allowed but are omitted here for simplicity, while the feature matrix $X\in\mathbb{R}^{n\times d}$ represents the node features. Here, $V$ and $E$ represent the sets of nodes and edges, respectively, with $|V|=n$ indicating the number of nodes. We use the notation $v_{1}\sim_{A}v_{2}$ to denote that $v_{1}$ and $v_{2}$ are neighboring nodes according to the adjacency $A$. Additionally, we define $[n]\coloneq\{1,2,\ldots n\}$, and $\mathcal{P}([n])$ as the power set of $[n]$.
Subgraph GNNs as graph products. In a recent work, [3] demonstrated that various types of update rules used by current Subgraph GNNs can be simulated by employing the Cartesian graph product between the original graph and another graph, and running standard message passing over that newly constructed product graph. Formally, the cartesian product of two graphs $G_{1}$ ($n_{1}$ nodes) and $G_{2}$ ($n_{2}$ nodes), denoted by $G_{1}\square G_{2}$, forms a graph with vertex set $V(G_{1})\times V(G_{2})$. Two vertices $(u_{1},u_{2})$ and $(v_{1},v_{2})$ are adjacent if either $u_{1}=v_{1}$ and $u_{2}$ is adjacent to $v_{2}$ in $G_{2}$, or $u_{2}=v_{2}$ and $u_{1}$ is adjacent to $v_{1}$ in $G_{1}$. We denote by $\mathcal{A}\in\mathbb{R}^{n_{1}\cdot n_{2}\times n_{1}\cdot n_{2}}$ and $\mathcal{X}\in\mathbb{R}^{n_{1}\cdot n_{2}\times d}$ the adjacency and node feature matrices of the product graph; in general, we use calligraphic letters to denote the adjacency and feature matrices of product graphs, while capital English letters are used for those of the original graphs. In particular, for the graph cartesian product, $G_{1}\square G_{2}$, the following holds:
@@
For a detailed definition of the cartesian product of graphs, please refer to Definition A.1. As a concrete example for the analogy between Subgraph GNNs and the Cartesian product of graphs, we refer to a result by [3], which states that the maximally expressive node-based Subgraph GNN architecture GNN-SSWL$+$ [39], can be simulated by an MPNN on the Cartesian product of the original graph with itself, denoted as $G\square G$. As we shall see, our framework utilizes a cartesian product of the original graph and a coarsened version of it, as illustrated in Figure 1 (right).
Equivariance. A function $L:U\rightarrow W$ is called equivariant if it commutes with the group action. More formally, given a group element, $g\in\mathbb{G}$, the function $L$ should satisfy $L(g\cdot v)=g\cdot L(v)$ for all $v\in U$ and $g\in\mathbb{G}$. $L$ is said to be invariant if $L(g\cdot v)=L(v)$.
@@ 4 Coarsening-based Subgraph GNN
Overview. This section introduces the Coarsening-based Subgraph GNN (CS-GNN) framework. The main idea is to select and process subgraphs in a principled and flexible manner through the following approach: (1) coarsen the original graph via a coarsening function, $\mathcal{T}$ – see Figure 1(left); (2) Obtain the product graph – Figure 1(right) defined by the combination of two adjacencies, $\mathcal{A}_{\mathcal{T}(G)}$ (red edges), $\mathcal{A}_{G}$ (grey edges), which arise from the graph Cartesian product operation (details follow); (3) leveraging the symmetry of this product graph to develop symmetry-based updates, described by $\mathcal{A}_{\text{Equiv}}$ (this part is not visualized in Figure 1). The general update of our suggested layer takes the following form 3,
@@
where the superscript t indicates the layer index. In what follows, we further elaborate on these three steps (in Sections 4.1, 4.1 and 4.2).
We note that each connectivity in Equation 2 is processed using a distinct MPNN, and after stacking of those layers, we apply a pooling layer [^4] For some of the theoretical analysis, this pooling operation is expressed as: $\rho(\mathcal{X}^{\mathtt{T}})=\texttt{MLP}^{\mathtt{T}}\left(\sum_{S}\left(\texttt{MLP}^{\mathtt{T}}\big{(}\sum_{v=1}^{n}\mathcal{X}^{\mathtt{T}}(S,v)\big{)}\right)\right)$ to obtain a graph representation; that is, $\rho(\mathcal{X}^{\mathtt{T}})=\texttt{MLP}^{\mathtt{T}}\Big{(}\sum_{S}\Big{(}\sum_{v=1}^{n}\mathcal{X}^{\mathtt{T}}(S,v)\Big{)}\Big{)}$; $\mathtt{T}$ denotes the final layer.
For more specific implementation details, we refer to Appendix F.
4.1 Construction of the coarse product graph
As mentioned before, a maximally expressive node-based Subgraph GNN can be realized via the Cartesian product of the original graph with itself $G\square G$. In this work, we extend this concept by allowing the left operand in the product to be the coarsened version of $G$, denoted as $\mathcal{T}(G)$, as defined next. This idea is illustrated in Figure 1.